Ramon took a collection of color titles from a box. Amelia took 13 titles from his collection and Kelko took half of those remaining. Ramon had 11 left . How many did he he start with?

Answer by ewatrrr(24785)

 
Hi
Let x represent the number of tiles Ramon took from the box
Amelia took 13 titles from his collection and Kelko took half of those remaining
Question states***Ramon had 11 left
x-13 = (x-13)/2 + 11 || Note: idea is: half = half
x/2 = 13/2 + 11
x = 13 + 22
x = 35